# From Semantics, Scene to Instance-awareness: Distilling Foundation Model for Open-vocabulary Situation Recognition

### v1

@@ 4.4. Qualitative results
The figure 4 presents a qualitative comparison of our proposed MIPD framework with ClipSitu (Roy et al., 2024) and CMD-SE (Lei et al., 2024c) on unseen (top) and rare (bottom) situations. In the unseen example, where the ground truth activity is Sitting, MIPD accurately predicts the activity along with all corresponding entities and their roles. In contrast, ClipSitu and CMD-SE misclassify the activity as Leaning and Slouching, respectively, and fail to identify several semantic roles correctly. In the rare example involving the activity Buckling, MIPD again aligns well with the ground truth, accurately detecting entities like Baby, Seatbelt, Car, and Man. In contrast, both ClipSitu and CMD-SE misidentify the activity and incorrectly label several semantic roles. These examples highlight MIPD’s ability to generalize beyond seen data, demonstrating its robustness in handling both unseen and rare grounded situation recognition.

### v2

@@ 4.4. Qualitative results
The figure 4 presents a qualitative comparison of our proposed MIPD framework with ClipSitu (Roy et al., 2024) and CMD-SE (Lei et al., 2024b) on unseen (top) and rare (bottom) situations. In the unseen example, where the ground truth activity is Sitting, MIPD accurately predicts the activity along with all corresponding entities and their roles. In contrast, ClipSitu and CMD-SE misclassify the activity as Leaning and Slouching, respectively, and fail to identify several semantic roles correctly. In the rare example involving the activity Buckling, MIPD again aligns well with the ground truth, accurately detecting entities like Baby, Seatbelt, Car, and Man. In contrast, both ClipSitu and CMD-SE misidentify the activity and incorrectly label several semantic roles. These examples highlight MIPD’s ability to generalize beyond seen data, demonstrating its robustness in handling both unseen and rare grounded situation recognition.